OpenStudy (seratul):

Anyone know an easy way to factor quadratic equations like these without graphing or quadratic formula?

OpenStudy (seratul):

\[14x^2+41x+15=0\]

OpenStudy (seratul):

@mathmate

OpenStudy (seratul):

@satellite73

OpenStudy (steve816):

Yeah, there is a method where you split the middle term.

OpenStudy (mathmate):

@Seratul Do you know how to factor integers? Like 24=2*2*2*3?

OpenStudy (seratul):

Yes.

OpenStudy (mathmate):

Good! The method work like the following: Let's say we have 15x^2+19x+6 and we'd like to factorize (without using quadratic formula).

OpenStudy (mathmate):

...where a=15, b=19, c=6 Then we have ac=15*6=90 And we look for two integers, m & n such that m*n=90, m+n=19 so far so good?

OpenStudy (seratul):

So far so good :)

OpenStudy (mathmate):

Can you find the two numbers m and n?

OpenStudy (seratul):

Oh, 10 and 9.

OpenStudy (mathmate):

Good, now, let me figure out the next step! lol

OpenStudy (seratul):

Lmao. Take your time. I appreciate you helping me btw.

OpenStudy (mathmate):

Then 9x and 10x would represent what you need to add together to get 19x.

OpenStudy (seratul):

Alright.

OpenStudy (mathmate):

In this case, it is relatively easy, because 9=3*3, and 10=5*2

OpenStudy (seratul):

And 5 times 3 is 15?

OpenStudy (mathmate):

yes, so it would be (5x + )(3x+ ) which you can fill easily.

OpenStudy (seratul):

How would I fill it?

OpenStudy (mathmate):

(5x + p )(3x+ q ) You need p*3x=9x, and 5x+q=10x

OpenStudy (seratul):

Why would q have a variable if we are trying to get a whole number of 6?

OpenStudy (mathmate):

Well, p*q \(should\) be 6 from what we have already done at the beginning. You can solve p and find q from pq=6 also.

OpenStudy (seratul):

Oh. I guess I can just do trail and error, so would it be (5x+3)(3x+2)=0?

OpenStudy (mathmate):

Exactly!

OpenStudy (seratul):

Thank you! Is it alright if I try to do 14x^2+41x+15=0 and just write my steps here?

OpenStudy (mathmate):

It's just I think there is an even easier way with less guessing, but please do and work on this one, using the same steps.

OpenStudy (mathmate):

So ac=14*15=210, and b=41 Did you find m and n?

OpenStudy (seratul):

Okay. 14*15=210 m*n=210 m+n=41 21+20=41 21=7,3 20=5,2,2 14= 7*2 So we can write (7x+q)(2x+p)=0

OpenStudy (seratul):

Oh, sorry. I went a little too quick.

OpenStudy (seratul):

Does this seem correct so far?

OpenStudy (mathmate):

You need to first find m and n. Your first try was 21+20, but the product is not 210. You are better off trying the correct product (210) first, and I'll show you how to get closer to the sum of 41.

OpenStudy (mathmate):

Can you give me any pair with m*n=210, don't worry about the sum.

OpenStudy (seratul):

Oh crap. Alright, 21 and 10 should work.

OpenStudy (mathmate):

Good, 21*10=210, first step satisfied. But 21+10=31....too small, right?

OpenStudy (seratul):

Unfortunately to small :(

OpenStudy (mathmate):

When the sum is too small, you \(stretch\) the range, get the big number bigger, and try a smaller than the small.

OpenStudy (mathmate):

Any tries?

OpenStudy (seratul):

I was gonna say 70 and 3, but that stretches a little too much.

OpenStudy (mathmate):

So you know that 35<m<70, and 3<n<10. Any more candidates?

OpenStudy (seratul):

5 and 42. Closer...

OpenStudy (mathmate):

*21<m<70 lol

OpenStudy (mathmate):

So you know which way to go?

OpenStudy (seratul):

Oh lol... Hmm, 35 and 6 :D

OpenStudy (mathmate):

Exactly! Not too bad, right?

OpenStudy (mathmate):

Next step.

OpenStudy (seratul):

It seems okay so far.

OpenStudy (seratul):

Okay, we break down 35 and 6. 35 - 7,5 6 - 3,2

OpenStudy (mathmate):

I'll show you an easier way, the method came back to me. Write (14x+35)(px+q)

OpenStudy (seratul):

May I ask where those numbers came from?

OpenStudy (mathmate):

So that we know 14 comes from a, 35 comes from m. (it could be 14 and n, but we'll try that later)

OpenStudy (mathmate):

We don't need to do the p, q, actually. We'll just write (14x+35). Do you see a common factor?

OpenStudy (seratul):

7

OpenStudy (mathmate):

Exactly, since the given expression does not have a common factor, so the 7 does not belong to the first factor (therefore the second). Now we have (2x+5)(7x+?) right?

OpenStudy (seratul):

Ah, yes.

OpenStudy (mathmate):

Can you finish it, before we try (14x+n)

OpenStudy (seratul):

It has to equal 15 so 3?

OpenStudy (seratul):

(2x+5)(7x+3)=0

OpenStudy (mathmate):

Exactly, so we just factorized it to be (2x+5)(7x+3), check (especially the first few times using this method). 2x*7x=14x^2, 2x(3)+(5)(7x)=6x+35x=41x, 5*3=15, yay!

OpenStudy (mathmate):

Now, what if we did (14x+6) can you do the rest?

OpenStudy (seratul):

Okay, so they have 2 in common. So (7x+3)(2x+5) Oh, its the same thing.

OpenStudy (mathmate):

So this shows that you can just choose any of m or n to start, and get the same result. Is everything clear?

OpenStudy (seratul):

Yes, thank god. Makes it so much clearer. Thank you very much. Do you have time for one more though? I want to try with a negative number.

OpenStudy (mathmate):

Sure, let's do it!

OpenStudy (seratul):

:D 45x^2-74x-55

OpenStudy (mathmate):

But remember to take care of the sign for m and n!

OpenStudy (mathmate):

I want to see you factor it! :)

OpenStudy (mathmate):

Do it step by step, don't jump!

OpenStudy (seratul):

Okay. Hopefully I don't mess up. 45*-55= -2475

OpenStudy (mathmate):

Excellent. Next step would be...

OpenStudy (seratul):

We need to figure out m and n

OpenStudy (mathmate):

Good! Keep going!

OpenStudy (mathmate):

Remember, try any product that gives -2475 and go from there. Also, since the sum is -74, you expect the larger number to be negative.

OpenStudy (seratul):

m*n=-2475 m+n=-74 25-99=-74

OpenStudy (mathmate):

Excellent! Keep the momentum there! Go, go ,go!

OpenStudy (seratul):

(45x+25x) = (9x+5)(5x+q)

OpenStudy (mathmate):

Now the last little step! Go, go, go!! You're almost there!!!!

OpenStudy (seratul):

5q=-55 q=-11 (9x+5)(5x-11)=0????

OpenStudy (mathmate):

Yep, but do a FOIL anyway!

OpenStudy (seratul):

OKAY!! 9x*5x-99x+25x-55 45x^2-74x-55=0!!!! :DDDD

OpenStudy (mathmate):

Well done! Cheer leaders dancing! :)

OpenStudy (seratul):

Lmao. Thanks for sticking with me for the past hour lol. I extremely appreciate it. I was never so happy to answer the problem!

OpenStudy (seratul):

You sir are amazing!

OpenStudy (mathmate):

No problem! I suggest you write up the method for future reference, or perhaps bookmark this post as well! If I forget (again), I'll ask you! lol

OpenStudy (seratul):

Of course. I'll probably be using this for the rest of algebra 2/trig.

OpenStudy (seratul):

Have a good day or night :)

OpenStudy (mathmate):

Good! Talk to you another time! :)